OpenStudy (anonymous):

Can you help me with this problem, please? 16/ -3 1/2

OpenStudy (anonymous):

Is this it? \(\Large\frac{16}{-3 \frac{1}{2}}\)

OpenStudy (anonymous):

Nope, it's another mixed fraction right?

OpenStudy (anonymous):

\[16-3 \frac{ 1 }{ 2 }\]

OpenStudy (anonymous):

Ah.

OpenStudy (anonymous):

What did you get this time justy?

OpenStudy (anonymous):

What's the Divide sign for..?

OpenStudy (anonymous):

After the 16.

OpenStudy (anonymous):

Oh I see, I didn't even see that haha...you may be right.

OpenStudy (anonymous):

Lol

OpenStudy (anonymous):

Lack of sleep makes me miss a lot of things, not very good.

OpenStudy (anonymous):

I guess even batman needs rest >.<

OpenStudy (anonymous):

no, iambatman, the problem is 16/ -3 1/2.

OpenStudy (anonymous):

Well once you do the mixed fraction use this rule, \[\huge \frac{ a }{ (\frac{ b }{ c }) } = \frac{ ac }{ b }\]

OpenStudy (anonymous):

LOL!

OpenStudy (anonymous):

Yeah I got that lol

OpenStudy (anonymous):

Dude, you should be asleep..how could you fight criminals at night if you don't SLEEP IN THE DAY?!

OpenStudy (anonymous):

I never sleep ;), anyways @Justy I see your problem is \[\huge\frac{16}{-3 \frac{1}{2}},~~~ -3 \frac{ 1 }{ 2 }=?\] Lets do this step by step.

OpenStudy (anonymous):

What's -3+1/2?

OpenStudy (anonymous):

im trying to solve the problem now. ill tell you soon. thanks (:

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Np :)

OpenStudy (anonymous):

is the answer 4.57?

OpenStudy (anonymous):

-4.57 yes

OpenStudy (anonymous):

Or -4 4/7

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

Np